OpenStudy (anonymous):

I forgot how to do this Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5).

OpenStudy (anonymous):

I know im suppose to plug (-2) and g(5) in though

OpenStudy (anonymous):

all youre doing is plugging and solving

OpenStudy (anonymous):

I dont know what to do after that

OpenStudy (anonymous):

Yeah.......

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5). F(-2) = 3 (-2) - 1 = -7 g(x) = -x + 6 g(5) = -5 + 6 = 1 I would say this is how you do it, but im not to sure. Sorry. Do you have choices?

OpenStudy (anonymous):

-6 6 8

OpenStudy (anonymous):

are the choices

OpenStudy (anonymous):

Well idk, dont take what i say as the right answer but i think i would go with -6... -7 + 1 = -6 (what i got for f(x) and g(x).

OpenStudy (anonymous):

Hope I kinda helped but im not that bright in math. Hopefully someone else can help you out better.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/525dd035e4b0ede9e44aaa04.. this person had the same question

OpenStudy (anonymous):

Did that help? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good :) turns out it was -6.. I seriously jus winged it but I seemed to be right.